OpenStudy (anonymous):

What number is less twice its reciprocal gives 23/5?

OpenStudy (radar):

X-2/X = 23/5 Multiply thru by 5 getting: 5X-10/X= 23 Multiply thru by X getting: \[5X ^{2}-10=23X\] Rearrange getting: \[5X ^{2}-23X-10=0\] Factor getting: (X-5)(5X+2)=0 Solve getting X=5, and X=-2/5